10.00 points The following materials standards have been established for a particular product: Standard quantity per unit of output Standard price 4.9 grams $12.00 per grams The following data pertain to operations concerning the product for the last month 3,800 grams Actual materials purchased Actual cost of materials purchased Actual materials used in production Actual output S 44,270 3,100 grams 570 units The direct materials purchases variance is computed when the materials are purchased Required a. What is the materials price variance for the month? (Input the amount as a positive value. Leave no cells blank - be certain to enter O wherever required. Indicate the effect of each variance by selecting F for favorable, U for unfavorable, and None for no effect (i.e., zero variance.) Materials price variance b. What is the materials quantity variance for the month? (Input the amount a as positive value. Leave no cells blank - be certain to enter O wherever required. Indicate the effect of each variance by selecting F for favorable, U for unfavorable, and None for no effect (i.e., zero variance.) Materials quantity variance
Answer #1

18)
a)Direct material price=(PQ*AP)-(PQ*SP)
PQ=Actual material
AP=Actual price
SP=standard price
=344270-(3800*12)=1330 and it is favourable
b)Direct material quantity=SP(AQ-Sq)  
SP=standard price
AQ= actual qunatity used
SQ=STANDARD QUANTITY ALLOWED
=(3100-(4.9*570))*12=3684 and it id unfavourble
19)
a)Direct labour rate=(AH*AR)-(AH*SR)  
AH=actual hours used
AR=actual rate per hour
SR=standard rate per hour
=125150-(9500*14)=7850 and it is favourable
b)Direct labour efficency=SR(AH-SH)  
SR=standard rate per hour'
AH=actual hours used
SH=standard hours allowed
=(9500-(1570*6))*14=1120 and it is unfavourable